# Electricity Price Forecasting: Bridging Linear Models, Neural Networks and Online Learning

### v1

@@ 3.3 Ensembling
$\hat{P}_{d,h}^{\text{ens}}=\sum_{k=1}^{K}w_{k,d,h}\hat{P}_{k,d,h}$
where $\hat{P}_{d,h}^{\text{ens}}$is the ensemble forecast at day $d$ and hour $h$ , $\hat{P}_{k,d,h}$ is the forecast of model $k$, $w_{k,d,h}$ is the corresponding weight, and $K$ is the number of models.
The simplest ensemble approach is to assign an equal weight to all models such that $w_{k,d,h}=\frac{1}{K}$. This weighting aggregation is employed, for example, by Random Forest ([breiman2001random]), where the forecasts of different trees (experts) are averaged. However, it is important to highlight that the strength of Random Forest lies not only in ensembling, but also in the construction of decorrelated trees and random feature selection [hastie2009elements, Ch. 15].
A more advanced ensemble method is the fully adaptive Bernstein Online Aggregation (BOA) method proposed by [wintenberger2017optimal]. The weight of this approach varies among experts and between time steps for an individual expert. It is expressed as follows
$w_{k,d,h}=\frac{\eta_{k,d-1,h}\,\exp\!\big(-\eta_{k,d-1,h}R_{k,d-1,h}\big)}{\sum_{j=1}^{K}\eta_{j,d-1,h}\,\exp\!\big(-\eta_{j,d-1,h}R_{j,d-1,h}\big)},$
where $\eta_{k,d-1,h}$ is the adaptive learning rate for expert $k$, and $R_{k,d-1,h}$ is the cumulative corrected regret of expert $k$, computed up to day d1d-1 for hour $h$. The inclusion of $R_{k,d-1,h}$ inside the exponential ensures that experts who performed poorly in the past (up to day d1d-1) receive a smaller weight in day $d$, while larger weights are assigned to the best-performing experts in the past. Moreover, unlike classical ensemble methods such as Exponentially Weighted Average (EWA) ([vovk1990aggregating]), fully adaptive BOA employed the second-order refinement that stabilises the learning process.
@@ 4.1 Learning Algorithm
The issue with this mechanism is that in each iteration, the model is refitted on the last $D$ observations with only a single additional observation compared to the previous iteration, which makes the procedure time-consuming. To reduce this computational burden, online learning is introduced such that instead of refitting all $D$ observations, the weights are updated in each iteration using the formerly estimated coefficients along the new observation. Online learning is applied in electricity price forecasting for both point and probabilistic predictions. For instance, [hirsch2024online] implements an algorithm that embeds online coordinate descent (OCD) or recursive least squares (RLS) into the Generalized Additive Models for Location, Scale and Shape (GAMLSS) framework to estimate regularized linear distributional models. Additionally, [yan2012online] employs an online/incremental Support Vector Regression (SVR) approach, which updates the model parameters as new data arrive, instead of retraining from scratch, to generate point forecasts of electricity prices.
@@
This project, on the other hand, implements a partial online learning using a novel non-parametric approach depicted in Figure 4. In the first iteration, the model is initialized randomly and fitted using 10 observations. The resulting weights are then employed to forecast observation 11. The optimal weights from the first iteration are used as initialization of the model in the second iteration, and only the most recent five observations are used to refit the model. Similarly, in iteration 3, the optimal weights from iteration 2 are carried forward, and the model is retrained on the latest five observations. This process continues till the final observation $T$ is forecasted. This approach is constructed on the assumption that the model requires a larger sample in the first iteration to learn the data structure, while in subsequent iterations, fewer observations suffice, since the carried-over weights already contain information from past data. The initial 10 observations are referred to as the initial window size, whereas the subsequent five observations used in each iteration define the update window size. The name partial comes from the use of a couple of observations in the updated iterations, not the new observation only, which is the standard in the online learning algorithm.
Figure 4: Visualisation of the online learning algorithm
@@
The window size visualization in Figure 4 serves only as a simple illustration of how the online learning procedure functions. However, there are other essential elements for the algorithm to work effectively, notably: the number of epochs, which determines the number of times the model sees the full dataset, and the learning rate, which controls the magnitude of parameter updates during training. The number of epochs in the initial iteration should generally be larger than the number of epochs in the updated windows. Since a higher number of epochs is associated with a larger number of weight updates, this ensures convergence toward a local minimum. However, for the subsequent iteration, the model starts in the neighborhood of the minimum thanks to initialization weights from the previous iteration, and then a small number of epochs is usually adequate to ensure convergence. A similar principle applies to the learning rate.
@@ 4.2 Study Design
The dataset is split into training, validation, and test sets, where the test data comprises the last two years of observations from 2023-01-16 to 2025-01-14, and the validation set includes the two prior years of observations. Training data, on the other hand, changes in every iteration. For instance, during the hyperparameter tuning phase, where the best hyperparameters are selected, the training set is the $D_{\text{init}}$ observations prior to 2021-01-16, which marks the first observation in the validation dataset. The second training set then comprises the $D_{\text{up}}$ observations preceding the second validation point, and so on, where $D_{\text{init}}$ is the window size of the initial window and $D_{\text{up}}$ is the window size of the updated window.
@@
Both $D_{\text{init}}$ and $D_{\text{up}}$ are hyperparameters to be tuned, along with the number of neurons in the hidden layer and initial and updated learning rate. In addition, the penalty terms $\lambda_{1}$ and $\lambda_{2}$ in the loss function (as defined in the following equation) are also subject to tuning:
$L(\bm{\theta})=\ell(\mathbf{P})+\lambda_{1}\|\bm{\theta}\|_{2}^{2}+\lambda_{2}\|\bm{w}^{(\mathrm{out})}\|_{1}$
@@
where $\ell(\mathbf{P})$ is L1 loss, $\|\bm{\theta}\|_{2}$ is the Euclidean norm of the weight vector $\bm{\theta}$, and $\|\bm{w}^{(\mathrm{out})}\|_{1}$ is norm of the output-layer weights. The use of this regularized loss function is necessary to avoid overfitting, which is particularly important given the large number of regressors. The last hyperparameter to consider for tuning is solely employed by the two OLS-based models. It is denoted as $\alpha$ and it determines the share of the OLS coefficient, $\bm{\hat{\beta}}_{\text{OLS}}$, to be used for initializing the skip connection of Reduced Linear with OLS and the MLP with Reduced Linear and OLS models. For example, if $\alpha=1$, the initialization is given directly by $\bm{\hat{\beta}}_{\text{OLS}}$, while if $\alpha=2$, then models will use $2\times\bm{\hat{\beta}}_{\text{OLS}}$ as initialisation.
@@
The Tree-structured Parzen Estimator (TPE) optimization method, proposed by [TPE], is employed to select the combination of optimal values of the above hyperparameters. TPE is a Bayesian optimization algorithm that selects candidate hyperparameters by modeling the distribution of good and bad trials from past evaluations to minimize the objective function. Unlike random search, which samples random values within the proposed interval in each trial, TPE employs previous trials to guide the search toward more promising regions of the hyperparameter space. We implement this hyperparameter optimization algorithm using the Python package Optuna ([optuna]), where we set the number of trials to 500 and set the mean absolute error as an objective function.
@@
On the other hand, there are other hyperparameters that we predetermine, such as the number of epochs, which we set to 60 for the initial iteration and 10 in the subsequent iterations, and the optimization algorithm, which we fix to Adaptive Moment Estimation (Adam) as it is simple and computationally efficient ([adam]).
@@
The resulting model for class $i$ will be denoted Name (BOA). For instance, the ensemble form resulting from the Reduced Linear architecture will be referred to as Reduced Linear (BOA). In addition to these seven class-specific ensembles, we also develop a general ensemble model, referred to as BOA all. This ensemble is built from the 500 models with the lowest MAE across all model classes. The same forward selection strategy is applied to identify the best combination of ten models. Note that all models, including individual and ensemble variants, are selected based on their performance on the validation dataset.
@@ 4.3 Benchmarks and Evaluation
In order to check the strengths of our models, two state-of-the-art models are selected as benchmarks. The first benchmark is Lasso Estimated AutoRegressive (LEAR), which is a high-dimensional linear benchmark for electricity price forecasting. LEAR coefficients are estimated using L1 regularization (LASSO) to enforce sparsity and prevent overfitting ([lear]). The second benchmark is the Deep Neural Network (DNN), which is a multilayer perceptron with input, output, and two hidden layers. DNN employs ADAM as an optimization algorithm for the weights and the Tree-structured Parzen Estimator algorithm to tune the hyperparameters ([dnn]).
@@
For each of the above accuracy metrics, the model with the smallest value of the corresponding metric is considered the best-performing model. However, there is no guarantee that this apparent superiority is statistically significant and could simply result from random chance. Hence, along with the accuracy metrics, the Diebold-Mariano (DM) test is employed, where the DM test is a statistical test that compares the forecast accuracy of two forecasting models ([dm]). Let $\hat{P}_{d,h}^{(A)}$ and $\hat{P}_{d,h}^{(B)}$ be the generated forecast at day $d$ on hour $h$ using models $A$ and $B$, respectively. The forecast error of model $A$ is $\hat{e}_{d}^{(A)}=\left(\sum_{h=0}^{23}\hat{P}_{d,h}^{(A)}-\sum_{h=0}^{23}{P}_{d,h}^{(A)}\right)$ and forecast error of model $B$ is $\hat{e}_{d}^{(B)}=\left(\sum_{h=0}^{23}\hat{P}_{d,h}^{(B)}-\sum_{h=0}^{23}{P}_{d,h}^{(B)}\right)$.
The null hypothesis to compare the forecast accuracy of $A$ and $B$ is denotes as $E\left(\left|\hat{e}_{d}^{(A)}\right|-\left|\hat{e}_{d}^{(B)}\right|\right)=0,$ where $E(.)$ is the expectation. The null hypothesis implies that there is no difference in forecast accuracy between $A$ and $B$. The alternative hypothesis is expressed as: $E\left(\left|\hat{e}_{d}^{(A)}\right|-\left|\hat{e}_{d}^{(B)}\right|\right)<0,$ which suggests that the forecasts of model $A$ are statistically more accurate than those of $B$.

### v2

@@ 3.3 Ensembling
$\hat{P}_{d,h}^{\text{ens}}=\sum_{k=1}^{K}w_{k,d-1,h}\hat{P}_{k,d,h}$
where $\hat{P}_{d,h}^{\text{ens}}$is the ensemble forecast at day $d$ and hour $h$ , $\hat{P}_{k,d,h}$ is the forecast of model $k$, $w_{k,d-1,h}$ is the corresponding weight, and $K$ is the number of models.
The simplest ensemble approach is to assign an equal weight to all models such that $w_{k,d-1,h}=\frac{1}{K}$. This weighting aggregation is employed, for example, by Random Forest ([breiman2001random]), where the forecasts of different trees (experts) are averaged. However, the strength of Random Forest lies not only in ensembling, but also in the construction of decorrelated trees and random feature selection [hastie2009elements, Ch. 15].
A more advanced ensemble method is the fully adaptive Bernstein Online Aggregation (BOA) method proposed by [wintenberger2017optimal]. The weight of this approach varies among experts and between time steps for an individual expert. The weights are computed as follows, using the implementation of the fully adaptive BOA presented in [berrisch2023crps]:
$r_{k,d,h}$ $=\ell^{\prime}(P_{d,h},\hat{P}_{d,h}^{ens})\left(\hat{P}_{k,d,h}-\hat{P}_{d,h}^{ens}\right)$
[…]
$V_{k,d,h}$ $=V_{k,d-1,h}+r_{k,d,h}^{2}$
$\eta_{k,d,h}$ $=\min\left(\sqrt{\frac{-\log(w_{k,0,h})}{V_{k,d,h}}},\frac{1}{2E_{k,d,h}}\right)$
$R_{k,d,h}$ $=R_{k,d-1,h}+\frac{1}{2}\Bigl(r_{k,d,h}\bigl(1+\eta_{k,d,h}r_{k,d,h}\bigr)\Bigr)+E_{k,d,h}\mathbf{1}\{2\eta_{k,d-1,h}r_{k,d,h}>1\}$
$w_{k,d,h}$ $=\frac{w_{k,0,h}\,\eta_{k,d,h}\,\exp(-\eta_{k,d,h}R_{k,d,h})}{\sum_{j=1}^{K}w_{j,0,h}\,\eta_{j,d,h}\,\exp(-\eta_{j,d,h}R_{j,d,h})},$
[…]
@@ 4.1 Learning Algorithm
The issue with this mechanism is that in each training window, the model is refitted on the last $D$ observations with only a single additional observation compared to the previous training window, which makes the procedure time-consuming. To reduce this computational burden, online learning is introduced such that instead of refitting all $D$ observations, the weights are updated in each training window using the formerly estimated coefficients along the new observation. Online learning is applied in electricity price forecasting for both point and probabilistic predictions. For instance, [hirsch2024online] implements an algorithm that embeds online coordinate descent (OCD) or recursive least squares (RLS) into the Generalized Additive Models for Location, Scale and Shape (GAMLSS) framework to estimate regularized linear distributional models. Additionally, [yan2012online] employs an online/incremental Support Vector Regression (SVR) approach, which updates the model parameters as new data arrive, instead of retraining from scratch, to generate point forecasts of electricity prices.
@@
This project, on the other hand, implements a partial online learning using a novel approach depicted in Figure 4. In the first training window, the model is initialized randomly and fitted using 10 observations. The resulting weights are then employed to forecast observation 11. The optimal weights from the first training window are used as initialization of the model in the second training window, and only the most recent five observations are used to refit the model. Similarly, in training window 3, the optimal weights from training window 2 are carried forward, and the model is retrained on the latest five observations. This process continues till the final observation $T$ is forecasted. This approach is constructed on the assumption that the model requires a larger sample in the first training window (referred to as the initial training window) to learn the data structure, while in subsequent training rounds (referred to as updated training windows) , fewer observations suffice, since the carried-over weights already contain information from past data. The initial 10 observations are referred to as the initial window size, $D_{\text{init}}$, whereas the subsequent five observations used in each updated training window define the update window size $D_{\text{up}}$. $D_{\text{init}}=10$ and $D_{\text{up}}=5$ serve only as a simple illustration of how the online learning procedure functions. In practice, both $D_{\text{init}}$ and $D_{\text{up}}$ are two hyperparameters that are tuned using the validation dataset.
Figure 4: Visualisation of the online learning algorithm for $D_{\text{init}}=10$ and $D_{\text{up}}=5$
@@
The window size is not the only component for the proposed partial online learning to work effectively. Other elements are equally essential, as they also take different values in the initial training window than in the updated training window, notably: the number of epochs, which determines the number of times the model sees the full dataset, and the learning rate, which controls the magnitude of parameter updates during training. The number of epochs in the initial training window should generally be larger than the number of epochs in the updated windows. Since a higher number of epochs is associated with a larger number of weight updates, this ensures convergence toward a local minimum. However, for the subsequent training window, the model starts in the neighborhood of the minimum thanks to initialization weights from the previous training window, and then a small number of epochs is usually adequate to ensure convergence. A similar principle applies to the learning rate and the penalty terms in the loss function defined in Equation 5. Note that since the window size is not the only component of the proposed online learning framework, there may be cases in which $D_{init}$ is smaller than $D_{up}$ (a configuration that is uncommon in standard online learning), yet the resulting approach remains faster than training without the proposed framework, due to the influence of the other hyperparameters.
The transfer of model parameters from one training round to the next, referred to as warm starting, is widely used in convex optimization problems. For example, [he2014practical] employ warm starting to incrementally update logistic regression models. In contrast, warm starting is rarely used in deep neural networks ([ash2020warm]). However, it is important to emphasize that the novelty of the proposed partial online learning framework does not stem solely from the use of warm starting. Rather, it lies in combining warm starting with window-specific hyperparameter configurations, where the initial and updated windows use hyperparameters tailored to their respective stages. Moreover, the term partial comes from the use of a couple of observations in the updated training rounds, $D_{up}>1$, not the new observation only, $D_{up}=1$, which is the standard in the online learning algorithm. Although setting $D_{up}=1$ is possible under the proposed learning approach, it is typically suboptimal in practice.
@@ 4.2 Study Design
The dataset is split into training, validation, and test sets, where the test data comprises the last two years of observations from 2023-01-16 to 2025-01-14, and the validation set includes the two prior years of observations (2021-01-16 to 2023-01-15). Training data, on the other hand, changes in every training window. For instance, during the hyperparameter tuning phase, the training set is the $D_{\text{init}}$ observations prior to 2021-01-16, which marks the first observation in the validation dataset. The second training set then comprises the $D_{\text{up}}$ observations preceding the second validation point, and so on.
Each of the models passes through two main phases: hyperparameter tuning and evaluation. During the first stage, the best hyperparameters (e.g, number of neurons) are selected based on the performance on the validation dataset. In the later phase, the model with the selected hyperparameters is assessed on the test dataset, which remains entirely unseen during training and tuning. During the two stages, the models are recalibrated daily, such that the network weights are updated in each training window. While the daily refit of models is common (or even advisable) during the evaluation stage, it is not typically performed in the tuning phase. For instance, the benchmark model, DNN, is fitted only once during tuning. However, due to the unique structure of the proposed online learning, where the value of most hyperparameters varies depending on the initial and updated training window, we follow a daily recalibration even in the tuning phase. This approach allows us to identify hyperparameters that are optimal for the corresponding training window.
@@
The hyperparameters to be tuned are $D_{\text{init}}$ and $D_{\text{up}}$, along with the number of neurons in the hidden layer and initial and updated learning rate. In addition, the penalty terms $\lambda_{1}$ and $\lambda_{2}$ in the loss function (as defined in the following equation) are also subject to tuning:
$L(\bm{\theta})=\ell(\mathbf{P})+\lambda_{1}\|\bm{\theta}\|_{2}^{2}+\lambda_{2}\|\bm{w}^{(\mathrm{out})}\|_{1}$ (5)
@@
where $\ell(\mathbf{P})$ is L1 loss, $\|\bm{\theta}\|_{2}$ is the Euclidean norm of the weight vector $\bm{\theta}$, and $\|\bm{w}^{(\mathrm{out})}\|_{1}$ is norm of the output-layer weights. The use of this regularized loss function is necessary to avoid overfitting, which is particularly important given the large number of regressors. The last hyperparameter to consider for tuning is solely employed by the two OLS-based models. It is denoted as $\alpha$ and it determines the share of the OLS coefficient, $\bm{\hat{\beta}}_{\text{OLS}}$, to be used for initializing the skip connection of RLin with OLS and the MLP with RLin and OLS models. For example, if $\alpha=1$, the initialization is given directly by $\bm{\hat{\beta}}_{\text{OLS}}$, while if $\alpha=2$, then models will use $2\times\bm{\hat{\beta}}_{\text{OLS}}$ as initialisation.
@@
The Tree-structured Parzen Estimator (TPE) optimization method, proposed by [TPE], is employed to select the combination of optimal values of the above hyperparameters. TPE is a Bayesian optimization algorithm that selects candidate hyperparameters by modeling the distribution of good and bad trials from past evaluations to minimize the objective function. Unlike random search, which samples random values within the proposed interval in each trial, TPE employs previous trials to guide the search toward more promising regions of the hyperparameter space. We implement this hyperparameter optimization algorithm using the Python package Optuna ([optuna]), where we set the number of trials to 500 and set the mean absolute error as an objective function. The search space of the hyperparameters is illustrated in Table 4 in the Appendix. The use of 500 trials corresponds to evaluating 500 candidate sets of hyperparameters, aiming for a balance between achieving high forecast accuracy and keeping a moderate computational cost.
@@
On the other hand, there are other hyperparameters that we predetermine, such as the number of epochs, which we set to 60 for the initial training round and 10 in the subsequent training rounds, and the optimization algorithm, which we fix to Adaptive Moment Estimation (Adam) as it is simple and computationally efficient ([adam]).
@@
The resulting model for class $i$ will be denoted [Class Name] (BOA). For instance, the ensemble resulting from the RLin architecture will be referred to as RLin (BOA). In addition to these seven class-specific ensembles, we also develop a general ensemble model, referred to as BOA all. This ensemble is built from the 500 models with the lowest MAE across all model classes. The same forward selection strategy is applied to identify the best combination of ten models.
This forward selection strategy is not guaranteed to be optimal, since, for example, a model that performs badly when selecting $\mathcal{M}_{j}^{i}$ may become highly informative when combined with a later selected model $\mathcal{M}_{k}^{i}$, outperforming the ensemble that includes $\mathcal{M}_{j}^{i}$, for $1\leq j<k\leq 10$. However, this approach is computationally efficient, as it evaluates only 4456 ensemble combinations, compared to the $\binom{500}{10}$ possible combinations required by an exhaustive search. Note that all models, including individual and ensemble variants, are selected based on their performance on the validation dataset.
@@ 4.3 Benchmarks and Evaluation
In order to check the strengths of our models, two state-of-the-art models are selected as benchmarks. The first benchmark is Lasso Estimated AutoRegressive (LEAR), which is a high-dimensional linear benchmark for electricity price forecasting. LEAR coefficients are estimated using L1 regularization (LASSO) to enforce sparsity and prevent overfitting ([lear]). The second benchmark is the Deep Neural Network (DNN), which is a multilayer perceptron with input, output, and two hidden layers. DNN employs ADAM as an optimization algorithm for the weights and TPE algorithm to tune the hyperparameters ([dnn]), where the number of trials are set to 500.
@@
  • commodity prices ($\text{NGas}_{d-2}$, $\text{Oil}_{d-2}$, $\text{Coal}_{d-2}$, $\text{EUA}_{d-2}$). Note that those fuel-related prices were not included in the LEAR and DNN discussed in [a]. They are added here to offer a fair comparison between the benchmark and proposed models by employing a similar set of features.
For DNN, the hyperparameter optimization is performed over the following search space: the number of hidden neurons, activation function, weight initialization scheme, random seed, dropout rate, learning rate, scaling strategies, optional L1 regularization, and binary selection of regressors.
In contrast, the developed models tune a reduced set of hyperparameters, such that the only common hyperparameters with DNN are the number of neurons, the learning rate, and L1 regularization, which is systematically applied rather than treated as optional. Other tuned DNN hyperparameters are either predetermined (e.g., activation function) in the proposed models or not part of the modeling framework (e.g., dropout mechanism). The random seed and weight initialization are also not considered in the developed models, as their influence is limited to the first training window, which is considered minimal.
During the evaluation stage, both LEAR and DNN are recalibrated daily, and their forecast performance is evaluated using the test data. LEAR does not go through the hyperparameter tuning stage as it has a single hyperparameter, namely the Lasso regularization parameter, which is recalibrated daily during the evaluation stage. This approach enhances LEAR’s predictive accuracy but also increases the training time relative to a standard Lasso model with a fixed penalty determined earlier. DNN hyperparameters, on the other hand, are selected using the validation data, where the network is trained only once and not recalibrated daily.
It is worth emphasizing that throughout this study, the benchmark models employ the exact same test datasets as the proposed models. However, the validation datasets differ: the proposed models use the explicitly specified validation dataset in the beginning of Section 4.2, whereas the benchmark model, DNN, use the 42 weeks preceding the test data. Additionally, within this study, one year is defined as 365 days, except when discussing the window size of the benchmark models, where one year is defined as 364 days to remain consistent with the code provided in the epftoolbox library.
@@
For each of the above accuracy metrics, the model with the smallest value of the corresponding metric is considered the best-performing model. However, there is no guarantee that this apparent superiority is statistically significant and could simply result from random chance. Hence, along with the accuracy metrics, the Diebold-Mariano (DM) test is employed, where the DM test is a statistical test that compares the forecast accuracy of two forecasting models ([dm]). Let $\hat{P}_{d,h}^{(A)}$ and $\hat{P}_{d,h}^{(B)}$ be the generated forecast at day $d$ on hour $h$ using models $A$ and $B$, respectively. The forecast error of model $A$ is $\hat{e}_{d}^{(A)}=\left(\sum_{h=0}^{23}\hat{P}_{d,h}^{(A)}-\sum_{h=0}^{23}{P}_{d,h}\right)$ and forecast error of model $B$ is $\hat{e}_{d}^{(B)}=\left(\sum_{h=0}^{23}\hat{P}_{d,h}^{(B)}-\sum_{h=0}^{23}{P}_{d,h}\right)$.
The null hypothesis to compare the forecast accuracy of $A$ and $B$ is denotes as $E\left(\Delta_{d}\right)=0,$ where $E(.)$ is the expectation, $\Delta_{d}=\left|\hat{e}_{d}^{(A)}\right|-\left|\hat{e}_{d}^{(B)}\right|$ under L1 loss and $\Delta_{d}=\left(\hat{e}_{d}^{(A)}\right)^{2}-\left(\hat{e}_{d}^{(B)}\right)^{2}$ under L2 loss. The null hypothesis implies that there is no difference in forecast accuracy between $A$ and $B$ under the assumption that the loss differential, $\Delta_{d}$, is stationary with short memory. The alternative hypothesis is expressed as: $E\left(\Delta_{d}\right)<0,$ which suggests that the forecasts of model $A$ are statistically significantly more accurate than those of $B$.
The Diebold–Mariano test is implemented using the small-sample correction proposed by [harvey1997testing] to control for finite-sample effects, and the long-run variance of the test statistic is estimated using the heteroskedasticity- and autocorrelation-consistent approach of [newey1987] with a lag truncation parameter of 7 to account for weekly seasonality.